# FLARE: Fast Low-rank Attention Routing Engine

### v2

@@ Abstract
The quadratic complexity of self-attention limits its applicability and scalability on large unstructured meshes. We introduce Fast Low-rank Attention Routing Engine (FLARE), a linear complexity self-attention mechanism that routes attention through fixed-length latent sequences. Each attention head performs global communication among $N$ tokens by projecting the input sequence onto a fixed length latent sequence of $M\ll N$ tokens using learnable query tokens. Routing information through this bottleneck constrains the interaction to have rank at most $M$: the resulting operation is equivalent to an attention matrix of rank at most $M$, realized at $\mathcal{O}(NM)$ cost. This formulation allows users to flexibly adjust the effective rank of attention –and thus the balance between accuracy and efficiency– by tuning $M$ to their resource and performance requirements. FLARE not only scales to unprecedented problem sizes, but also delivers superior accuracy compared to state-of-the-art neural PDE surrogates across diverse benchmarks. We also introduce a new additive manufacturing dataset to spur further research. Our code is available at https://github.com/vpuri3/FLARE.py.
@@ 1 Introduction
High-fidelity simulations of physical systems are often too costly for multi-query applications, such as design optimization or uncertainty quantification. Machine learning offers a promising alternative via surrogate models that learn system dynamics from data, enabling fast approximations that accelerate experimentation and decision making.
Among machine learning (ML) architectures, transformers (Vaswani et al., 2017) have shown exceptional scalability and generalization capabilities in domains such as natural language processing (Devlin et al., 2019) and computer vision (Dosovitskiy et al., 2020). This has led to increased interest in adapting transformers for spatially distributed data such as point clouds and meshes in physical simulations. However, applying transformers directly to large-scale unstructured meshes introduces severe computational bottlenecks. The core challenge lies in the self-attention mechanism in which, for a sequence of $N$ tokens, each token attends to every other token in the sequence, resulting in $\mathcal{O}(N^{2})$ time and memory complexity. Although this global communication is key to the expressive power of transformers, enabling them to outperform graph neural networks (GNNs) that are inherently local (Yun et al., 2019), its computational footprint hinders its scalability.
In the context of learning partial differential equation (PDE) surrogates, i.e., regressing continuous fields over point clouds/meshes, each point in the 3D point cloud is treated as a token, with associated features that encode geometric and physical quantities such as coordinates, normals, and material properties.
Several works have proposed approximations to the full self-attention mechanism to mitigate the cost. Perceiver and PerceiverIO (Jaegle et al., 2021b; a) introduced cross-attention between the input sequence and a length $M$ learnable latent sequence. In this mechanism, the $M\times N$ attention weights are used to project the input sequence to the latent sequence, followed by self-attention blocks in the latent space. This decouples the length of the input sequence from the computational complexity by avoiding dense self-attention on the inputs. However, the latent bottleneck in PerceiverIO can limit accuracy as the model may discard fine-grained features if the number of latent tokens is too low.
In learning PDE surrogates, Transolver (Wu et al., 2024) and Latent Neural Operator (LNO) (Wang & Wang, 2024a) adopt a similar projection-unprojection scheme to map variable-length point clouds to fixed-length latent representations. Both models obtain projection weights by expanding the feature dimension, and apply self-attention on the resulting latent sequence. The differences between the three approaches are highlighted in Table 1. Notably, Perceiver and LNO only perform a single encoding and decoding step, whereas Transolver reapplies projection and unprojection in every transformer block. This enables Transolver to form deep models, where attention connections in downstream layers benefit from the rich embeddings and global context aggregated by previous layers. However, the attention projector layers proposed in Transolver and LNO cannot utilize existing GPU kernels for scaled dot-product attention (SDPA), a well-optimized algorithm in modern deep learning (Dao et al., 2022). This limits their scalability to large meshes.
Table 1: Comparison of FLARE with different latent attention schemes in PerceiverIO (Jaegle et al., 2021a), LNO (Wang & Wang, 2024a), and Transolver (Wu et al., 2024).
Criteria PerceiverIO Transolver LNO FLARE (ours)
Sequential encoding and decoding blocks
Parallel encodings & decodings
Weight sharing b/w encoding and decoding
Latent space self-attention
Compatible with fused attention kernels
  • *Physics Attention in Transolver uses same projection weights for all heads whereas PerceiverIO and FLARE use cross-attention projection where each head learns a distinct slice of the latent query tokens.
We propose Fast Low-rank Attention Routing Engine (FLARE), a simple yet effective mechanism designed to break the scalability barrier in PDE surrogate learning. FLARE is built on the argument that projecting input sequences onto shorter latent sequences, and then unprojecting to the original sequence length, is equivalent to constructing a low-rank form of attention with rank at most equal to the number of latent tokens (see Figure 1).
Furthermore, we argue that multiple simultaneous low-rank projections could collectively capture a full attention pattern. Unlike Transolver which shares projection weights across heads, or LNO which applies only a single projection, our design allocates a distinct slice of the latent tokens to each head resulting in distinct projection matrices for each head. This allows each head to learn independent attention relationships, opening up a key direction of scaling and exploration, wherein each head may specialize in distinct routing patterns.
To support this intuition, we draw on known properties of rank-deficient matrices, to develop a linear-time eigenanalysis procedure to study the learned communication matrices. Our spectral analysis reveals that attention heads operate in low-rank subspaces, validating our core design choice. Furthermore, we observe diversity in the eigen spectrum across heads, reinforcing the motivation for independent head-wise projections. These findings highlight the efficiency and expressiveness of our low-rank formulation. We summarize our main contributions below.
Figure 1: Schematic of a FLARE block. In FLARE, each head projects the input sequence with NN tokens to a fixed-length sequence of $M$ tokens via the cross-attention matrix $W_{\text{encode}}=\mathrm{softmax}(Q\cdot K^{T})$, and then projects back to the original length via the cross-attention matrix $W_{\text{decode}}=\mathrm{softmax}(K\cdot Q^{T})$. The overall operation is equivalent to token mixing on the input sequence with the rank-deficient matrix $\left(W_{\text{decode}}\cdot W_{\text{encode}}\right)$.
  • Linear complexity token mixing. FLARE is self-attention mechanism designed to learn on long sequences such as point clouds. By replacing full self-attention with low-rank projections and reconstructions, FLARE achieves linear complexity in the number of points.
  • Superior accuracy. Across multiple PDE benchmarks, FLARE achieves superior predictive accuracy compared to leading neural surrogate models, despite operating with fewer parameters, and at much lower computational complexity.
  • Scalability. FLARE is built entirely from standard fused attention primitives, ensuring high GPU utilization and ease of integration into existing transformer architectures. As such, FLARE enables end-to-end training on unstructured meshes with one million points (see Figure 2) without distributed computing (Luo et al., 2025) or memory offloading – the largest scale demonstrated for transformer-based PDE surrogates.
@@
  • Benchmark dataset of additive manufacturing simulations. We release a large-scale, high-resolution dataset for residual displacement prediction in metal additive manufacturing to encourage further research in scalable PDE surrogate modeling. The dataset is available at https://huggingface.co/datasets/vedantpuri/LPBF_FLARE.
@@ 2 Related work
Learning neural surrogates for PDEs is an exciting and rapidly growing direction in scientific machine learning. Neural operators (Li et al., 2020; Lu et al., 2021; Kovachki et al., 2023) have been proposed for learning mappings between infinite-dimensional input and output function spaces, enabling mesh-independent generalization. Extensions to neural operators that incorporate geometric priors and graph-based representations further improve performance, especially for problems defined on complex unstructured meshes (Li et al., 2023b; c).
Graph networks have also been widely explored as PDE surrogates (Pfaff et al., 2020; Elrefaie et al., 2024; Ferguson et al., 2025), leveraging their inherent ability to model local neighborhood interactions on meshes. Transformer-based architectures have emerged more recently as powerful PDE surrogates (Alkin et al., 2024; Cao, 2021; Li et al., 2022; 2023a; Hao et al., 2023), allowing global context aggregation and dynamic prediction of complex phenomena (Qian et al., 2025). Recent works (Alkin et al., 2024; Wu et al., 2024; Wang & Wang, 2024a) leverage latent space attentions for PDE modeling, achieving high accuracy with reduced computational cost. Building upon these advances, our FLARE method employs a linear complexity attention mechanism with a learnable latent bottleneck to further improve accuracy and efficiency on PDE surrogate modeling tasks.
@@
Several methods have been proposed to address the quadratic complexity of the attention mechanism. In (Wang et al., 2020), attention matrices are shown to be low-rank. Their proposed model, called Linformer, employs learned linear mappings to project the input key and value sequences to a shorter sequence length, resulting in a linear complexity form of attention. A drawback of this approach is that it cannot directly handle variable input sequence lengths. Other low-rank approaches include Reformer (Kitaev et al., 2020), which uses locality-sensitive hashing for efficient approximate attention, Nyströmformer (Xiong et al., 2021), which approximates self-attention using the Nyström method, and Scatterbrain (Chen et al., 2021), which combines sparse and low-rank methods for further gains. Another efficient attention method, called multi-head latent attention (Liu et al., 2024), jointly compresses key and value tokens to reduce the memory footprint of language models during inference. However, this does not help with the quadratic bottleneck in self-attention. Low-Rank Adaptation (LoRA) (Hu et al., 2022) also leverages low-rank projections, though primarily for efficient fine-tuning rather than architectural design or modeling capacity, as we explore in this work.
Figure 2: Time and memory requirements of different attention schemes. On an input sequence of one million tokens, FLARE (red) is over $200\times$ faster than vanilla attention, while consuming marginally more memory. All models are implemented with flash attention (Dao et al., 2022), and the memory upper bound on a single H100 80GB GPU is depicted with a dashed line. Note that the curves for FLARE are somewhat overlapping.
@@ 3.1 Preliminary: Multi-Head Self-Attention
Let $X\in\mathbb{R}^{N\times C}$ denote the input sequence of $N$ tokens with $C$ features each. The query, key, and value matrices $Q,K,V\in\mathbb{R}^{N\times C}$ are obtained by applying learned linear projections to $X$, we have
@@
$\left[Q_{1},\ldots,Q_{H}\right]=Q,\hskip 10.00002pt\left[K_{1},\ldots,K_{H}\right]=K,\hskip 10.00002pt\left[V_{1},\ldots,V_{H}\right]=V.$ (2)
@@
The greatest cost in multi-head self-attention is the call to SDPA which is $\mathcal{O}(N^{2})$ in time and memory complexity. This is because $Q_{h}\cdot K_{h}^{T}\in\mathbb{R}^{N\times N}$ requires $\mathcal{O}(N^{2})$ storage and softmax, matrix-vector product with $V_{h}$ takes $\mathcal{O}(N^{2})$ operations. Fortunately, GPU optimized multi-head implementations of SDPA are available in PyTorch (Paszke, 2019).
1import torch.nn.functional as F
2def flare_multihead_mixer(q, k, v):
3 # Args - q: [H, M, D], k, v: [B, H, N, D]
4 # Ret - y: [B, H, N, D]
5 z = F.scaled_dot_product_attention(q, k, v, scale=1.0)
6 y = F.scaled_dot_product_attention(k, q, z, scale=1.0)
7 return y
@@ 3.2 FLARE: Fast Low-rank Attention Routing Engine
Formally, we define a learnable query matrix $Q\in\mathbb{R}^{M\times C}$, where each row corresponds to a latent token. The key and value matrices, $K,V\in\mathbb{R}^{N\times C}$, are obtained by applying deep residual multi-layer perceptrons (MLPs) detailed in Appendix A to the input $X$. Compared to just a linear layer, these allow the model to learn higher-order feature interactions and deeper nonlinear transformations. Refer to Appendix F for ablation studies.
@@
where $Y_{h}\in\mathbb{R}^{N\times D}$ is the output for each head. Similar to multi-head self attention, the outputs from all heads are concatenated along the feature dimension and passed through a final linear projection to mix information across heads. As the query matrix has only $M$ tokens, the cost of SDPA calls in Eq. 5 and Eq. 6 is $\mathcal{O}(NM)$. PyTorch code for the multi-head implementation is presented in Figure 3. Note that we use a scaling factor $s=1$ instead of $\sqrt{D}$ in typical transformers (Vaswani et al., 2017)) in SDPA. This modification is explained in Appendix F.
@@
$W_{\text{encode},h}$ $=\mathrm{softmax}(Q_{h}\cdot K_{h}^{T})\in\mathbb{R}^{M\times N},\,\text{and}$ (8)
@@
as the dense global communication matrix with rank at most $M$. This low-rank structure, illustrated in Figure 1, enables efficient all-to-all communication without explicitly forming $W_{h}$; instead, $W_{\text{encode},h}$ and $W_{\text{decode},h}$ are applied sequentially, resulting in an overall cost of $\mathcal{O}(MN)$ per head.
@@
We defer to Appendix E a deeper analysis of how FLARE’s latent tokens enable gather–scatter communication (acting as selective pooling hubs and broadcasters), why the symmetry between encoding and decoding operators promotes stable information flow, and how fixing latent queries highlights a tradeoff between query dynamics and the need for expressive key/value projections. Together, these perspectives clarify the structural principles that underlie FLARE’s efficiency and expressivity.
@@
Here, ResMLP (He et al., 2016) denotes a residual MLP block detailed in Appendix A, and LayerNorm denotes layer normalization (Ba et al., 2016) operation. To summarize, a FLARE block consists of a token mixing operation via FLARE, a pointwise residual MLP, and layer normalization in pre-norm format (Xiong et al., 2020). Deep residual MLPs within the block enable complex, token-level feature transformations and improve accuracy.
@@
The overall architecture is given by $B$ sequential FLARE blocks sandwiched between an input projection and an output projection, which are detailed in Appendix A. Such a design enables the model to efficiently integrate local and global information across multiple layers, making it well suited for large-scale, high-dimensional data such as point clouds.

### v3

@@ Abstract
The quadratic complexity of self-attention limits the scalability of transformers on long sequences. We introduce Fast Low-rank Attention Routing Engine (FLARE), a token-mixing operator that realizes low-rank attention by routing information through a small set of latent tokens. Each layer induces an input-input token mixing matrix of rank at most $M$ via a minimal encode-decode factorization implemented using only two standard scaled dot-product attention (SDPA) calls. Because the dominant $\mathcal{O}(NM)$ computation is expressed purely in terms of standard SDPA, FLARE is compatible with fused attention kernels and avoids materializing $M\times N$ projection matrices. FLARE further assigns disjoint latent slices to each attention head, yielding a mixture of head-specific low-rank pathways. Empirically, FLARE scales to one-million-point unstructured meshes on a single GPU, achieves state-of-the-art accuracy on PDE surrogate benchmarks, and outperforms general-purpose efficient-attention methods on the Long Range Arena suite. We additionally release a large-scale additive manufacturing benchmark dataset. Our code is available at https://github.com/vpuri3/FLARE.py
Machine Learning, ICML
@@ 1 Introduction
High-fidelity simulations of physical systems are often too costly for multi-query applications such as design optimization and uncertainty quantification. Neural surrogate models offer a promising alternative by learning mappings from simulation inputs to outputs, enabling rapid evaluation once trained.
Transformers (Vaswani et al., 2017) have demonstrated strong scalability and generalization across domains including natural language processing (Devlin et al., 2019) and computer vision (Dosovitskiy et al., 2020). This success has motivated their adoption for spatially distributed scientific data such as point clouds and unstructured meshes, where each mesh point is treated as a token with geometric and physical features. However, standard self-attention scales as $\mathcal{O}(N^{2})$ in both time and memory for $N$ tokens, which becomes prohibitive for high-resolution meshes containing hundreds of thousands to millions of points.
A common strategy for reducing attention cost is to introduce a low-dimensional latent bottleneck of size $M\ll N$, reducing complexity to $\mathcal{O}(NM)$. Low-rank attention methods such as Linformer (Wang et al., 2020) assume that attention matrices are approximately low-rank and learn explicit $M\times N$ projection matrices. While effective at moderate sequence lengths, this approach requires $\mathcal{O}(NM)$ parameters per layer and implicitly assumes a fixed token ordering, making it impractical for very long or unordered sequences such as unstructured meshes. Latent-token architectures such as Perceiver and PerceiverIO (Jaegle et al., 2021b, a), Transolver Wu et al. (2024), and Latent Neural Operator (LNO) Wang and Wang (2024b) instead use attention projectors to map inputs to a fixed latent array, followed by deep latent-space self-attention. Notably, most latent-token methods introduce additional transformations inside the latent space, making the overall attention behavior a composition of projections, latent mixing, and nonlinear refinements rather than a single, interpretable operator.
Figure 1: FLARE block. Each head encodes NN input tokens into $M$ latent tokens with cross-attention $W_{\mathrm{enc}}=\mathrm{softmax}(Q\cdot K^{T})$ and decodes back with $W_{\mathrm{dec}}=\mathrm{softmax}(K\cdot Q^{T})$, inducing an explicit rank-$\leq M$ input-space mixing operator $W=W_{\mathrm{dec}}\cdot W_{\mathrm{enc}}$. Unlike PerceiverIO (Jaegle et al., 2021a), Transolver (Wu et al., 2024), and LNO (Wang and Wang, 2024b), FLARE uses no latent self-attention and executes both steps with fused SDPA kernels (Dao et al., 2022) (see Figure 6 for a summary).
Figure 2: Time and memory requirements of different attention schemes. On an input sequence of one million tokens, FLARE (red) is over $200\times$ faster than vanilla attention, while consuming marginally more memory. All models are implemented with FlashAttention (Dao et al., 2022), and the memory upper bound on a single H100 80GB GPU is depicted with a dashed line. Note that the curves for FLARE are somewhat overlapping. A detailed analysis is presented in Appendix G.
We introduce Fast Low-rank Attention Routing Engine (FLARE), a token-mixing layer that realizes a flexible low-rank self-attention operator by routing information through a small set of latent tokens. Concretely, each layer induces an input-input token mixing matrix of rank at most $M$ via an encode-decode factorization implemented only with two standard scaled dot-product attention (SDPA) calls; no additional token-mixing is performed. This formulation makes the induced operator directly analyzable (e.g., via spectra) and keeps the dominant $\mathcal{O}(NM)$ computation fully SDPA-compatible, enabling fused attention kernels without materializing $M\times N$ projection weights.
FLARE operationalizes this idea with a minimal architecture: each block performs exactly two cross-attention operations (encode and decode). Relative to Linformer (Wang et al., 2020), FLARE parameterizes low-rank token mixing without learned $M\times N$ projection matrices or a fixed token ordering. Relative to latent-token architectures such as PerceiverIO (Jaegle et al., 2021a), Transolver (Wu et al., 2024), and LNO (Wang and Wang, 2024b), which use latents as a computational workspace and introduce additional latent token mixing, FLARE uses latents only for routing: all token mixing is realized by a single encode-decode factorization on the input tokens, eliminating latent-space self-attention and yielding a mathematically explicit attention operator. Counterintuitively, we find through systematic ablation that this architectural simplification (removing latent self-attention entirely) consistently improves accuracy while reducing parameters compared to prior latent-attention methods.
FLARE further assigns disjoint latent slices to each attention head, yielding a mixture of head-specific low-rank operators rather than a single shared bottleneck. As a result, each head implements an independent low-rank projection–reconstruction pathway, and global communication arises from aggregating multiple such operators. We analyze these operators using a linear-time eigenanalysis method and observe diverse eigenvalue decay patterns across heads. A controlled shared-latent ablation collapses these spectra and consistently degrades accuracy, indicating that independent latent routing is an important contributor to expressivity. Together, these results support the interpretation that FLARE benefits from combining complementary low-rank pathways rather than relying on a single shared projection.
Taken together, the architectural features of FLARE, (i) explicit encode-decode low-rank factorization, (ii) elimination of latent-space self-attention, and (iii) head-wise independent latent routing, emerge from systematic empirical study as key principles for efficient, scalable attention. This minimal design enables FLARE to achieve state-of-the-art performance on PDE surrogate benchmarks, scale to one-million-point unstructured meshes on a single GPU, and outperform general-purpose efficient-attention methods, all while using fewer parameters than existing approaches.
Figure 1 illustrates a single FLARE block. We summarize our main contributions below.
  • Flexible low-rank attention via latent routing. We introduce a self-attention formulation in which routing through latent tokens induces an implicit rank-$\leq M$ attention operator on the input sequence. FLARE eliminates all latent-space self-attention and expresses attention as a single encode-decode factorization, yielding a mathematically transparent operator.
  • Head-wise independent low-rank pathways. By assigning each attention head its own latent slice, FLARE learns a mixture of parallel low-rank operators; targeted ablations show this head-wise independence is critical for accuracy.
  • Extreme scalability with strong empirical performance. FLARE trains end-to-end on one-million-point unstructured meshes on a single GPU, achieves state-of-the-art PDE surrogate accuracy, and outperforms efficient-attention baselines on Long Range Arena.
@@
  • Benchmark dataset for additive manufacturing. We release a large-scale thermomechanical additive manufacturing dataset to support research on scalable scientific surrogates. The dataset is available at https://huggingface.co/datasets/vedantpuri/LPBF_FLARE.
@@ 2 Related work
Neural operators (Li et al., 2020; Lu et al., 2021; Kovachki et al., 2023) learn mappings between function spaces and enable mesh-independent generalization. Extensions incorporating geometric priors and graph-based representations improve performance on unstructured meshes (Li et al., 2023b, c). Transformer-based architectures have emerged more recently as powerful PDE surrogates (Alkin et al., 2024; Cao, 2021; Li et al., 2022, 2023a; Hao et al., 2023), allowing global context aggregation. Recent works (Alkin et al., 2024; Wu et al., 2024; Luo et al., 2025; Wang and Wang, 2024b) leverage latent space attentions for PDE modeling, achieving high accuracy with reduced computational cost. Building upon these advances, FLARE presents a linear-complexity attention operator tailored to extreme-resolution unstructured domains.
@@
Efficient attention methods reduce cost by constraining attention structure (e.g., rank, sparsity, or locality), yielding restricted subsets of full self-attention with bounded expressivity. Linformer (Wang et al., 2020) reduces attention complexity by learning explicit low-rank projections, but requires $\mathcal{O}(NM)$ parameters per layer. PerceiverIO (Jaegle et al., 2021a) and related models introduce latent tokens to reduce computational cost, using latent-space self-attention as a computational workspace to aggregate and transform global information. In this design, interactions between input tokens are mediated through a sequence of latent-space transformations rather than defined by a single, explicit operator on the input tokens. Transolver (Wu et al., 2024; Luo et al., 2025) and LNO (Wang and Wang, 2024b) adapt this latent projection-unprojection paradigm for PDE surrogate modeling. FLARE differs from these approaches by using latent tokens solely to induce an explicit low-rank operator, eliminating latent self-attention entirely and enabling a simpler, more analyzable attention formulation.
@@ 3.1 Preliminary: Multi-Head Self-Attention
Let $X\in\mathbb{R}^{N\times C}$ denote the input sequence of $N$ tokens with $C$ features each. The query, key, and value matrices $Q,K,V\in\mathbb{R}^{N\times C}$ are obtained by applying learned linear projections to $X$:
@@
$\left[Q_{1},\ldots,Q_{H}\right]=Q,\,\left[K_{1},\ldots,K_{H}\right]=K,\,\left[V_{1},\ldots,V_{H}\right]=V.$ (2)
@@
The greatest cost in multi-head self-attention is the call to SDPA which is $\mathcal{O}(N^{2})$ in time and memory complexity. This is because $Q_{h}\cdot K_{h}^{T}\in\mathbb{R}^{N\times N}$ requires $\mathcal{O}(N^{2})$ storage and softmax, matrix-vector product with $V_{h}$ takes $\mathcal{O}(N^{2})$ operations. Fortunately, GPU-optimized multi-head implementations of SDPA are available in PyTorch (Paszke, 2019).
@@ 3.2 FLARE: Fast Low-rank Attention Routing Engine
Formally, we define a learnable query matrix $Q\in\mathbb{R}^{M\times C}$, where each row corresponds to a latent token. The key and value matrices, $K,V\in\mathbb{R}^{N\times C}$, are obtained by applying deep residual multi-layer perceptrons (MLPs) detailed in Appendix B to the input $X$. Compared to just a linear layer, these allow the model to learn higher-order feature interactions and deeper nonlinear transformations. Refer to Appendix G for ablation studies.
@@
where $Y_{h}\in\mathbb{R}^{N\times D}$ is the output for each head. Similar to multi-head self-attention, the outputs from all heads are concatenated along the feature dimension and passed through a final linear projection to mix information across heads. As the query matrix has only $M$ tokens, the cost of SDPA calls in Eq. 5 and Eq. 6 is $\mathcal{O}(NM)$. PyTorch code for the multi-head implementation is presented in Figure 3. Note that we use a scaling factor $s=1$ instead of $\sqrt{D}$ in typical transformers (Vaswani et al., 2017) in SDPA. This modification is explained in Appendix G.
from torch.nn.functional import scaled_dot_product_attention as SDPA
def flare_multihead_mixer(Q, K, V):
# Args - Q: [H M D], K, V: [B H N D]
# Ret - Y: [B H N D]
Z = SDPA(Q, K, V, scale=1.0)
Y = SDPA(K, Q, Z, scale=1.0)
return Y
@@
$W_{\text{encode},h}$ $=\mathrm{softmax}(Q_{h}\cdot K_{h}^{T})\in\mathbb{R}^{M\times N},\quad\text{and}$ (8)
@@
as the dense global communication matrix with rank at most $M$. This low-rank structure, illustrated in Figure 1, enables efficient all-to-all communication without explicitly forming $W_{h}$; instead, $W_{\text{encode},h}$ and $W_{\text{decode},h}$ are applied sequentially, at an overall cost of $\mathcal{O}(MN)$ per head.
Model Elasticity Darcy Airfoil Pipe DrivAerML-40k LPBF
Vanilla Transformer (Vaswani et al., 2017) 5.37 660k 4.38 660k 6.28 660k $\sim$ $\sim$ $\sim$
PerceiverIO (Jaegle et al., 2021a) 28.0 1.87m 20.6 1.87m 7.65 1.87m 6.90 1.87m 248 1.87m 23.1 1.87m
GNOT (Hao et al., 2023) 13.3 4.87m 16.9 4.90m 103 4.90m 5.89 4.90m 115 4.87m 24.3 4.87m
LNO (Wang and Wang, 2024b) 9.25 1.83m 7.64 762k 17.8 762k 8.10 762k 146 762k 24.7 762k
Transolver w/o conv (Wu et al., 2024) 6.40 713k 18.6 713k 8.24 713k 4.87 713k 70.5 713k 20.4 713k
[…]
FLARE (ours) 3.38 592k 5.10 691k 4.28 691k 2.85 625k 60.8 691k 18.5 625k
@@
FLARE’s latent tokens implement a structured gather-scatter communication pattern: encoding pools input values into $M$ compact global descriptors, then decoding selectively broadcasts these back along the same pathways, yielding a low-rank routing operator that avoids explicit all-to-all interactions. This mechanism relies on two design choices discussed in Appendix F: (i) symmetric encode-decode operators promote stable information flow without redundant parameters, and (ii) fixed input-independent latent queries simplify the low-rank structure and improve efficiency, shifting expressive capacity to deep residual key/value projections that recover representational power.
@@
Here, ResMLP (He et al., 2016) denotes a residual MLP block detailed in Appendix B, and LayerNorm denotes layer normalization (Ba et al., 2016) operation. To summarize, a FLARE block consists of a token mixing operation via FLARE, a pointwise residual MLP, and layer normalization in pre-norm format (Xiong et al., 2020). Deep residual MLPs within the block enable complex, token-level feature transformations and improve accuracy.
@@
The overall architecture is given by $B$ sequential FLARE blocks sandwiched between an input projection and an output projection, which are detailed in Appendix B. Such a design enables the model to efficiently integrate local and global information across multiple layers, making it well suited for large-scale, high-dimensional data such as point clouds.
(a) Elasticity
(b) Darcy
(c) Pipe
Figure 4: Qualitative results for FLARE on the (a) Elasticity, (b) Darcy, and (c) Pipe datasets. We show the ground truth, the model prediction, and the corresponding error (Ground Truth - Prediction).